Find a1 of arithmetic sequence from given information.
a7=-45 a15=-77

Answer by AnlytcPhil(1806)

Find a1 of arithmetic sequence from given information.
a7 = -45 a15 = -77
Use the formula:
an = a1 + (n-1)d
a7 = a1 + (7-1)d
-45 = a1 + 6d
a15 = a1 + (15-1)d
-77 = a1 + 14d
So you have this system of equations:
-45 = a1 + 6d
-77 = a1 + 14d
Can you solve that system of equations?
If not post again.
a1 = -21, d = -4
Edwin
AnlytcPhil@aol.com